# Charged analogues of a singularity-free anisotropic compact star under linear $f(Q)$- action

### v3

@@ 3 Mathematical formalism: $f(Q)$ gravity and Einstein-Maxwell field equations
In this paper, we investigating a novel generalised approach to charged compact star modeling in the framework of the $f(Q)$ theory of gravity. Within the realm of gravitational theories, the concept of static spherically symmetric spacetime serves as a foundational assumption, offering valuable insights into various aspects of astrophysical phenomena. Keeping this in mind, we have considered a static spherically symmetric spacetime for the analysis of the present paper. Now, in the static spherically symmetric curvature co-ordinate space $(t,r,\theta,\phi)$, the line element is expressed as:
@@
Using Eq. (15) in Eq. (9), we obtain the form of non-metricity scalar as:
@@
where, $u_{i}$ represents the four velocity related to the anisotropic perfect fluid and $v_{i}$ is the unit space-like vector in the radial direction. These parameters obey the relations $u^{i}u_{i}=-1$ and $v^{i}v_{i}=1$. Moreover, $\rho,~{}p_{r},~{}p_{t}$ are the energy density and radial and tangential pressures, respectively, and are directed towards $u_{i}$ and $v_{i}$. The electromagnetic energy-momentum tensor takes the form:
@@
where, $q(r)$ is the total charge contained within the sphere of radius $r$ and is written as:
@@
For this study, we have considered a particular form of charge as follows [60]:
@@
where, $q_{0}$ is the intensity of charge and $q_{0}=0$ describes a neutral uncharged case.
@@
where, $\alpha_{0}$ has the dimension of $Km^{-2}$ and $\alpha_{1}$ is dimensionless. Now, using Eqs. (24, 25, 26, 27, 28), the field equations in the modified $f(Q)$ theory of gravity reduce to an exact set of formulations in the following form:
@@
Since, the components of gravitational field equations depend on the metric potentials, $\lambda(r)$ and $\nu(r)$, we must remember a subtlety, in choosing the forms of $\lambda(r)$ and $\nu(r)$, as proposed by Lake [61] which states that $\lambda(r)$ must be finite and constant at the centre $(r=0)$ and must hold the following properties: $\lambda^{\prime}(0)=0$ and $\lambda^{\prime\prime}(0)>0$. On the other hand, near the core, $\nu(r)$ must have the form $\nu=1+\mathcal{O}(r^{2})$, and it must increase towards the stellar boundary. This behaviour of the metric potentials, $\lambda(r)$ and $\nu(r)$ guarantees a monotonically decreasing energy density and pressure profile, with the maximum values concentrated in the core and diminishing steadily towards the surface. This characteristic is crucial for a stable and physically realistic stellar system. To fulfil the specified constraints in the present model, we opt for the Krori-Barua (KB) metric ansatz [62] in the form $\lambda=Ar^{2}$ and $\nu=Br^{2}+C$. Now, plugging the metric potentials along with Eq. (23) in the set of equations Eqs. (29), (30) and (31), we obtain the non-zero components of the field equations as:
@@
Now, the total active gravitational mass contained within the sphere of radius $R$ is,
@@ 4 Boundary conditions
In the framework of modified theories of gravity, the equations of motion as prescribed by GR are modified. Therefore, the associated boundary conditions applicable to any stellar system in GR require suitable modifications [63]. Now, in the 4-dimensional spacetime manifold $(\Omega)$, there exist two distinct regions, viz., the interior and exterior space-times. These regions are demarcated by a 3-dimensional hypersurface $(\Sigma)$ defined by the induced metric $h_{ij}$ in the $X^{i}$ coordinate system, with indices excluding the direction orthogonal to $\Sigma$. We express the projection tensors and the normal vector from $\Omega$ and $\Sigma$ in the form, $e^{a}_{i}=\frac{\partial x^{a}}{\partial X^{i}}$ and $n_{a}=\epsilon\delta_{a}\ell$, where, the affine connection in the perpendicular direction to $\Sigma$ is denoted by $\ell$ and for the respective timelike, null and spacelike geodesics, $\epsilon$ takes the values $-1,~{}0$ or $1$. In light of the above composition, one considers $n^{a}e^{i}_{a}=0$. Now, the induced metric $h_{ij}$ and extrinsic curvature tensor $K_{ij}$ at the hypersurface are written in the form:
@@
$ds^{2}=-\Bigg{(}1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\Bigg{)}~{}dt^{2}+\frac{dr^{2}}{\Bigg{(}1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\Bigg{)}}+r^{2}(d\theta^{2}+sin^{2}\theta d\phi^{2}),$ (38)
where $M=m(r=R)$ and $Q=q(r=R)$ are the total mass and charge of the stellar configuration respectively. Interestingly, in the parametric space of spherically symmetric space-time, $K_{ij}^{\pm}$ contains only two components viz., $K_{00}$ and $K_{\theta\theta}=K_{\phi\phi}sin^{2}\theta$, where the (+)(+)( + ) and ()(-)( - ) signs are for the exterior and interior space-time respectively. Using the KB ansatz and Eq. (38), we obtain the following components:
@@
$K_{\theta\theta}^{+}=\frac{r}{\sqrt{\frac{r^{2}}{Q^{2}-2Mr+r^{2}}}}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}K_{\theta\theta}^{-}=re^{-Ar^{2}}.$ (40)
Another necessary condition is the nullification of radial pressure at the boundary, i.e.,
@@
$A=\frac{ln\Big{[}\frac{R^{2}}{R^{2}-2MR+Q^{2}}\Big{]}}{2R^{2}},$ (42)
$B=\frac{2q_{0}^{2}R^{6}-\alpha_{0}R^{4}-2\alpha_{1}Q^{2}+4\alpha_{1}MR}{8\alpha_{1}R^{2}(Q^{2}-2MR+R^{2})},$ (43)
@@ 5 Physical characteristics of the proposed model
Assessing the physical viability of the model is crucial to ensure that the compact star is physically realistic. A stellar model must comply with the regularity and reality conditions persistent in the context of relativistic stellar modeling to qualify as a compact star. In this section, we have considered the compact star PSR J0740+6620 with a mass of $2.072~{}M_{\odot}$ and radius of $12.39~{}Km$ [73] to study the impact of charge on various stellar parameters and to test the physical viability of the present theoretical model within the framework of modified $f(Q)$ gravity.Notably, following the study of Maurya et al. [68], we have considered $\alpha_{0}=10^{-46}~{}Km^{-2}$ along with $\alpha_{1}=-0.5$.
@@ 5.1 Thermodynamic properties: energy density, radial and tangential pressures and pressure anisotropy
To ensure a realistic stellar model, the energy density and pressure profiles must have a monotonically decreasing nature. Moreover, for a physically acceptable solution, the condition of tangential pressure states that it must be positive throughout the interior of the compact star. The anisotropy in pressure, must be zero at the centre, i.e., $\Delta=0$, which reflects that $p_{r}=p_{t}$ at the stellar core. These claims are supported by Figs. 1, 2, 3 and 4.
@@
Figure 1: Radial variation of energy density $(\rho)$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@
Figure 2: Radial variation of radial pressure $(p_{r})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@
Figure 3: Radial variation of tangential pressure $(p_{t})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@
Figure 4: Radial variation of pressure anisotropy $(\Delta\times 10^{6})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@
Figure 5: Radial variation of energy density gradient $(\frac{d\rho}{dr}\times 10^{6})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@
Figure 6: Radial variation of radial pressure gradient $(\frac{dp_{r}}{dr}\times 10^{6})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@
Figure 7: Radial variation of tangential pressure gradient $(\frac{dp_{t}}{dr}\times 10^{6})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@
From Figs. 1, 2 and 3, it is noted here that, the energy density $(\rho)$, radial $(p_{r})$ and tangential pressures $(p_{t})$ attain a maximum value at the stellar core, and as the charge increases, these physical parameters decrease. The maximisation of the core density and pressure profiles is ensured by $\rho^{\prime}(0)=0$, $p^{\prime}_{r}(0)=0$ and $p^{\prime}_{t}(0)=0$, while the nature of the gradients is negative for all $0<r\leq R$. These behaviours of the gradients are shown in Figs. (5), (6) and (7). Moreover, Fig. (4) shows that $\Delta>0$ throughout the interior of the star, which denotes a repulsive nature pointing towards a more massive structure. In light of the above physical analysis, we can say that the thermodynamic properties are well satisfied in this model, which points towards a physically acceptable stellar model in stable equilibrium.
@@ 5.2 Maximum mass and radius from TOV equation
In this section, we numerically solve the TOV equations [69, 70] to obtain the maximum mass and radius for a charged compact star. It should be noted that, the solution of the TOV equation is entirely based upon the choice of a particular EoS. Hence, to resort to a suitable EoS, we adopt the method of maximising the radial sound velocity $(v_{r}^{2}=\frac{dp_{r}}{d\rho}=1)$, to the extreme end of causality, at the centre and the method of curve fitting from the data set of $\rho$ and $p_{r}$. We tabulate the EoS for increasing charge intensity $(q_{0})$ in Table 1.
@@
Using Table 1, we have numerically solved TOV equations and the results are plotted in Fig. 8. The obtained maximum mass and radius are tabulated in Table 2.
Figure 8: Mass-radius plot for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@
Charge intensity $(q_{0})$ Maximum mass $(M_{\odot})$ Radius (Km)
0.0002 2.838 12.0039
0.0004 2.849 12.0348
0.0006 2.869 12.0822
From Table 2, it is evident that the increase in charge intensity $(q_{0})$ results in an increase in the maximum mass in the present scenario, which denotes the transition of the EoS from a softer to a stiffer nature. Interestingly, in this section, we have predicted the radii of some recently observed pulsars and GW events and evaluated their physical characteristic parameters, viz., central density $(\rho_{c})$, surface density $(\rho_{s})$ and central pressure $(p_{c})$, and they are tabulated in Table 3.
@@
PSR J0740+6620 [73] $2.072^{+0.067}_{-0.066}$ $12.39^{-1.30}_{-0.98}$ 0.0006 11.82 $0.41\times 10^{15}$ $0.23\times 10^{15}$ $0.52\times 10^{35}$

### v4

@@ 3 Mathematical formalism: $f(Q)$ gravity and Einstein-Maxwell field equations
In this paper, we investigate a novel generalised approach to charged compact star modeling in the framework of the $f(Q)$ theory of gravity. Within the realm of gravitational theories, the concept of static spherically symmetric spacetime serves as a foundational assumption, offering valuable insights into various aspects of astrophysical phenomena. Keeping this in mind, a static spherically and symmetric spacetime has been considered for the analysis of the present paper. Now, in the static spherically symmetric curvature co-ordinate space $(t,r,\theta,\phi)$, the line element is expressed as:
@@
Using Eq. (15) in Eq. (9), we obtain the expression of non-metricity scalar as:
@@
where, $u_{i}$ represents the four velocity related to the anisotropic perfect fluid and $v_{i}$ represent the space-like unit vector along the radial direction. These parameters obey the relations $u^{i}u_{i}=-1$ and $v^{i}v_{i}=1$. The terms $\rho,~{}p_{r},~{}p_{t}$ in Eq. (17) represent respectively the energy density and pressures along the radial and tangential directions associated with the fluid. The electromagnetic energy-momentum tensor takes the form:
@@
where, $q(r)$ in Eq. (20) represents the total charge associated with the fluid sphere of radius $r$ and is written as:
@@
To examine the impact of electric charge on the structure of the present model, it is essential to define the charge distribution within the star. We assume that the electric charge varies according to a specific function, dependent on the radial coordinate. In the present analysis, we have considered a particular form of charge as follows [59]:
@@
where, $q_{0}$ is the intensity of charge having dimension of $Km^{-2}$ and $q_{0}=0$ describes a neutral uncharged case.
@@
where, $\alpha_{0}$ has the dimension of $Km^{-2}$ and $\alpha_{1}$ is dimensionless. Now, using Eqs. (24), (25), (26), (27) and (28), the field equations in the modified $f(Q)$ theory of gravity reduce to an exact set of formulations in the following form:
@@
Choice of metric potentials $\lambda(r)$ and $\nu(r)$ are very crucial as the components of gravitational field equations depend on them. Lake [61] proposed that $\lambda(r)$ should be finite and constant at the stellar centre $(r=0)$ as well as must hold the following properties: $\lambda^{\prime}(0)=0$ and $\lambda^{\prime\prime}(0)>0$. On the other hand, near the core, $\nu(r)$ must have the form $\nu=1+\mathcal{O}(r^{2})$, and it must increase towards the stellar boundary. Such behaviour of the metric potentials, $\lambda(r)$ and $\nu(r)$ guarantees a monotonically decreasing energy density and pressure profile. This characteristic is crucial for a stable and physically realistic stellar system. To fulfil the specified constraints in the present model, we opt for the Krori-Barua (KB) metric ansatz [62] in the form $\lambda=Ar^{2}$ and $\nu=Br^{2}+C$. Now, plugging the metric potentials along with Eq. (23) in the set of equations Eqs. (29), (30) and (31), we obtain the non-zero components of the field equations as:
@@
Now, the total active gravitational mass in presence of net charge contained within a spherical volume of radius $R$ is,
@@ 4 Boundary conditions
In the context of modified theories of gravity, the form of equations of motion as prescribed by GR are modified. Therefore, the associated boundary conditions applicable to any stellar system in GR require suitable modifications [63]. Now, in the 4-dimensional spacetime manifold $(\Omega)$, there exist two distinct regions, viz., the interior and exterior space-times. These regions are demarcated by a 3-dimensional hypersurface $(\Sigma)$ defined by the induced metric $h_{ij}$ in the $X^{i}$ coordinate system, with indices excluding the direction orthogonal to $\Sigma$. We express the projection tensors and the normal vector from $\Omega$ and $\Sigma$ in the form, $e^{a}_{i}=\frac{\partial x^{a}}{\partial X^{i}}$ and $n_{a}=\epsilon\delta_{a}\ell$, where, the affine connection in the perpendicular direction to $\Sigma$ is denoted by $\ell$ and for the respective timelike, null and spacelike geodesics, $\epsilon$ takes the values $-1,~{}0$ or $1$. In light of the above composition, one considers $n^{a}e^{i}_{a}=0$. Now, the induced metric $h_{ij}$ and the extrinsic curvature tensor $K_{ij}$ at the hypersurface are written in the form:
@@
$ds^{2}=-\Bigg{(}1-\frac{2M}{r}+\frac{Q_{c}^{2}}{r^{2}}\Bigg{)}~{}dt^{2}+\Bigg{(}1-\frac{2M}{r}+\frac{Q_{c}^{2}}{r^{2}}\Bigg{)}^{-1}dr^{2}+r^{2}(d\theta^{2}+sin^{2}\theta d\phi^{2}),$ (38)
where $M=$ the total mass and $Q_{c}=$ the total charge of the stellar configuration of radius $R$ respectively. Interestingly, in the parametric space of spherically symmetric space-time, $K_{ij}^{\pm}$ contains only two components viz., $K_{00}$ and $K_{\theta\theta}=K_{\phi\phi}sin^{2}\theta$, where the (+)(+)( + ) and ()(-)( - ) signs are for the exterior and interior space-time respectively. Using the KB ansatz and Eq. (38), we obtain the following components:
@@
$K_{\theta\theta}^{+}=\frac{r}{\sqrt{\frac{r^{2}}{Q_{c}^{2}-2Mr+r^{2}}}}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}K_{\theta\theta}^{-}=re^{-Ar^{2}}.$ (40)
Another necessary condition is that, at the surface of the star, radial pressure vanishes, i.e.,
@@
$A=\frac{ln\Big{[}\frac{R^{2}}{R^{2}-2MR+Q_{c}^{2}}\Big{]}}{2R^{2}},$ (42)
$B=\frac{2q_{0}^{2}R^{6}-\alpha_{0}R^{4}-2\alpha_{1}Q_{c}^{2}+4\alpha_{1}MR}{8\alpha_{1}R^{2}(Q_{c}^{2}-2MR+R^{2})},$ (43)
@@ 5 Physical characteristics of the proposed model
It is very crucial that any stellar model should be physically viable to ensure that the physical parameters associated with the star, obtained using the model, must be physically realistic. A stellar model must comply with the regularity and reality conditions persistent in the context of relativistic stellar modeling to qualify as a compact star. In this section, we have considered the compact star PSR J0740+6620 of estimated mass $2.072~{}M_{\odot}$ and radius $12.39~{}Km$ [73] to study the impact of charge on various stellar parameters and to test the physical viability of the present theoretical model within the framework of modified $f(Q)$ gravity. We have noted that the choice of charge intensity $(q_{0})$ in the present model, in the range of $0.0002-0.0009~{}Km^{2}$ yields a viable stellar model and notably, following the study of Maurya et al. [68], we have considered $\alpha_{0}=10^{-46}~{}Km^{-2}$ along with $\alpha_{1}=-0.5$ as arbitrary choices.
@@ 5.1 Thermodynamic properties: energy density, radial and tangential pressures and pressure anisotropy
To ensure a physically viable stellar model, the properties of energy density and pressures must have a nature of monotonically decreasing behaviour with r. Moreover, for a physically acceptable solution, the value of tangential pressure must be positive at all points interior of the compact star. The anisotropy in pressure, must be zero at the centre, i.e., $\Delta=0$, which reflects that $p_{r}=p_{t}$ at the stellar core. These claims are supported by Figs. 1, 2, 3 and 4.
@@
Figure 1: Radial variation of energy density $(\rho)$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
@@
Figure 2: Radial variation of radial pressure $(p_{r})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
@@
Figure 3: Radial variation of tangential pressure $(p_{t})$ for different charge intensities $(q_{0})$. Solid, dashed and dot-dashed lines correspond to $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
@@
Figure 4: Radial variation of pressure anisotropy $(\Delta\times 10^{-6})$ for different charge intensities $(q_{0})$. Solid, dashed and dot-dashed lines correspond to $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
@@
Figure 5: Radial variation of energy density gradient $(\frac{d\rho}{dr}\times 10^{-6})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines correspond to $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
@@
Figure 6: Radial variation of radial pressure gradient $(\frac{dp_{r}}{dr}\times 10^{-6})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
@@
Figure 7: Radial variation of tangential pressure gradient $(\frac{dp_{t}}{dr}\times 10^{-6})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines correspond to $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
@@
From Figs. 1, 2 and 3, it is noted here that the physical parameters such as energy density $(\rho)$ and pressures $(p_{r}$ and $p_{t})$ attain a maximum value at the stellar core, and as the charge increases, these physical parameters decrease. This may be attributed to the fact that as charge increases within a compact star, the resulting electrostatic repulsion counteracts the inward gravitational pull, leading to a drop in both energy density and pressure profiles. This effect arises because the outward force generated by the accumulated charge reduces the compression needed to sustain high densities and pressures, especially in the inner regions. Consequently, the overall matter distribution adjusts, balancing the electric repulsion with gravitational forces, resulting in lower density and pressure profiles as shown in Figs. 1, 2 and 3. Moreover, the maximisation of the core density and pressure profiles is ensured by $\rho^{\prime}(0)=0$, $p^{\prime}_{r}(0)=0$ and $p^{\prime}_{t}(0)=0$ and the nature of the gradients which is negative for all $0<r\leq R$. These behaviours of the gradients are shown in the Figs. 5, 6 and 7. Additionally, from Figs. 5, 6 and 7, we note that as the charge intensity $(q_{0})$ increases the profiles of the gradients increase. Moreover, Fig. (4) shows that $\Delta>0$ throughout the interior of the star and with increasing charge intensity $(q_{0})$, $\Delta$ decreases. Based on this behaviour, we may note that with increasing $q_{0}$, the electric field introduces an additional outward force which may help to equalise the pressure distribution by partially compensating for the intense inward gravitational pull, resulting in decreased pressure anisotropy. Furthermore, the positiveness of $\Delta$ denotes a repulsive anisotropic nature pointing towards a more massive structure. In light of the above physical analysis, we can say that the thermodynamic properties are well satisfied in this model, which points towards a physically acceptable stellar model in stable equilibrium.
@@ 5.2 Maximum mass and radius from TOV equation
In this section, we numerically solve the TOV equations [69, 70] to obtain the maximum mass of a star as well as its radius for a charged compact star. It should be noted that, the solution of the TOV equation is entirely based upon the choice of a particular EoS. Hence, to resort to a suitable EoS, we adopt the method of maximising the radial sound velocity $(v_{r}^{2}=\frac{dp_{r}}{d\rho}=1)$, to the extreme end of causality, at the centre and the method of curve fitting from the data set of $\rho$ and $p_{r}$. We tabulate the EoS for increasing charge intensity $(q_{0})$ in Table 1.
@@
0.0009 $p_{r}=0.908106~{}\rho-0.000352706$
Using Table 1, we have numerically solved TOV equations and the results are plotted in Fig. 8. Maximum mass and corresponding radius are evaluated from Fig. 8 and are tabulated in Table 2 for different charge intensities $(q_{0})$.
Figure 8: Mass-radius plot for different charge intensities $(q_{0})$. The solid, dashed, dot-dashed and large-dashed lines represent $q_{0}=0.0002,~{}0.0004,~{}0.0006$ and $0.0009~{}Km^{-2}$ respectively.
@@
Charge intensity $(q_{0})$ $(Km^{-2})$ Maximum mass $(M_{\odot})$ Radius (Km)
0.0002 2.84 12.00
0.0004 2.85 12.03
0.0006 2.87 12.08
0.0009 2.92 12.20
It is noted that maximum mass and radius both increase with charge intensity $(q_{0})$ in this model, which denotes the transition of the EoS from a softer to a stiffer nature. We have also predicted the radii of some pulsars and GW events observed recently and evaluated their characteristic of physical parameters, viz., central density $(\rho_{c})$, surface density $(\rho_{s})$ and central pressure $(p_{c})$, and they are tabulated in Table 3.
@@
PSR J0740+6620 [73] $2.072^{+0.067}_{-0.066}$ $12.39^{+1.30}_{-0.98}$ 0.0006 11.82 $0.41\times 10^{15}$ $0.23\times 10^{15}$ $0.52\times 10^{35}$